# The SOFIA Massive (SOMA) Radio Survey. II. Radio Emission from High-Luminosity Protostars

### v1

@@ 3.1.4 IRAS 16562-3959
Source 1 is the main detection in the region, and in our ATCA observations, this source is seen at all frequencies (5.5, 9, 17 and 22.8 GHz) and exhibits a morphology with a slight elongation in the NE-SW direction. SOMA II was able to resolve this source into two sources (G305B1 and G305B2) using high-spatial-resolution Gemini observations, finding morphologies that are also aligned in the NE-SW direction. They suggested that both sources are part of an outflow cavity. Our spectral index estimates show that the ionized material associated with this region is of non-thermal nature, suggesting the existence of a very fast jet in the region.
@@
Source 4 is a faint compact eastern source only seen at the lowest frequencies (5.5 and 9 GHz), and we report a limit on its spectral index (see Table 16). On the other hand, source 3 is seen at all frequencies and presents an elongated morphology in the SE-NW direction that appears as a double peak at the longest contour levels shown in Figure 1. This source may be part of the HII region in Walsh et al. (1999), but in our observations we report it as an independent source. We report a negative spectral index ( -0.6) suggesting non-thermal emission.
Since sources 1, 2, and 4 are aligned in the same SW-NW direction, exhibit relatively compact morphologies and negative spectral indices, we conclude that the most likely nature of these independent sources is that they are knots from a non-thermal jet. Following our convention for sources with jet-like morphology, we measured the flux densities at an intermediate scale that encompasses the central source and the jet knots (Sources 1, 2, and 4). At this scale, we estimate an upper limit on the spectral index of , which is indicative of a non-thermal jet in the region.
@@
CM1 | 19:23:06.87 | +14.20.17.3 | 0.8(0.1) | IR | No |
CM2A | 19:23:06.66 | +14.20.11.6 | 1.6(0.2) | IR | Yesa |
CM2B | 19:23:06.62 | +14.20.12.1 | 0.7 | IR | Yesa |
CM2C | 19:23:06.58 | +14.20.11.9 | 0.0(0.1) | IR | Yesa |
CM3 | 19:23:06.71 | +14.19.59.0 | 0.5 | IR | Yesb |
@@
Cyganowski et al. (2011) suggests that CM1 is a single ionizing star of spectral type B0V based on optically thin free–free emission and the ionizing photon flux. Our 1.3 cm flux density values are one order of magnitude less than that reported by Cyganowski et al. (2011); Towner et al. (2017) with VLA observations and resolutions of ′′ (2.63 and 1.62 mJy). This difference is likely due to the different resolutions.
CM2: The MIR peak observed in SOMA II is associated with source CM2, previously reported as an unresolved, 4 detection at 0.027 mJy/beam rms at 3.6 cm (Cyganowski et al., 2011). Additionally, Cyganowski et al. (2011) estimated a spectral index limit of 0.2 at a resolution of ′′. Using our VLA data, we are able to resolve source CM2 into three components, which we have labeled as CM2A, CM2B and CM2C, with all sources detected at both 1.3 and 6 cm observations.
@@
CM3: From the contour maps in Cyganowski et al. (2011); Towner et al. (2017), source CM3 shows MIPS 24 m emission and SOMA II also detected this source in the SOFIA IR maps, but not in the NIR. This source has no previous radio detections at 1.3 and 3.6 cm. In our observations we are able to detect this source, but only at 6 cm (C-band).
CM3 exhibits a compact morphology and we report a relative flat spectral index with an upper limit of 0.5, indicative of thermal emission. To our knowledge, we are detecting this source for the first time in radio frequencies and its naming follows the convention of the main region, G49.27-0.34.
@@
G339.88-1.26 (or IRAS 16484-4603) is a massive protostar located at 2.1 kpc (Krishnan et al., 2015), that has been previously observed at a variety of wavelengths. De Buizer et al. (2002) resolved the central MIR emission into three peaks (1A, 1B and 1C) and radio continuum observations have identified the central source is associated with a radio jet and the other two sources are associated with radio outflow lobes (Ellingsen et al., 1996; Purser et al., 2016), label as C, NE and SW. Using ALMA data, Zhang et al. (2019b) reported two molecular outflows towards this region positioned at the source C. One of this outflows is located in the northeast-southwest orientation and exhibits the same angle as the ionized outflow reported by Purser et al. (2016).
@@
For our analysis we used S. Purser’s ATCA observations at 5.5, 9.0, 17.0, and 22.8 GHz from Purser et al. (2016). In our analysis we consider the same three sources NE, C, and SW reported by Purser et al. (2016).
Source C presents a compact morphology at higher frequencies, but shows a slight extension at the lower frequencies, particularly at 5.5 GHz. We report a spectral index of 0.9. Meanwhile, sources NE and SW have spectral indices of 0.4 and 0.8, respectively. Both sources exhibit relatively compact morphologies, especially source SW.
As mentioned by Purser et al. (2016), source G339.88-1.26 shows characteristics consistent with an ionized radio jet driven by the central source C and two lobes of non-thermal emission. Given this jet-like morphology of the region we also measure the flux densities using an intermediate scale that encompass the three sources. We report a spectral index of 0.1 for the intermediate scale suggesting that the non-thermal emission from the lobes is more prevalent at this scale.
@@ 3.2 Radio SEDs
In Figure 5 we present the E-SEDs, i.e., radio + IR SEDs, for our nine protostellar sources. The dashed lines correspond to the best fit to the data using a power law of the form , where is the spectral index at the various scales: ”SOMA”; “Intermediate”; and “Inner”, as described above. The spectral index was calculated using the flux density at the central frequencies from the images, so is calculated over a wide frequency range (20 GHz). The uncertainty in the spectral index was calculated with a Monte Carlo simulation that bootstrapped the flux density uncertainties. We estimated an upper limit in the spectral index for non-detections at higher frequencies using a value of of 3.
It is important to mention that at higher frequencies the free-free emission is likely contaminated by dust emission (see Brogan et al. (2016)), therefore adding some additional systematic uncertainty to the measurement of the free-free fluxes.
@@ 4.1 Multiplicity and new detections
To make sure we are not missing any UCHII regions, we estimate the physical properties from the 5 GHz sensitivity limit (rms) using equation 1 and 3 from Kurtz et al. (1994) and equation A.2.3 from Panagia & Walmsley (1978). These equations assume spherical symmetry and optically thin emission from a uniform-density plasma with . The results are listed in Table 18, where column 1 is the region name, column 2 is the logarithm of the Lyman continuum flux () required for ionization. With the and the results from Table II in Panagia (1973), we estimate the spectral type of the ionizing star (listed in column 3), further assuming that a single ZAMS star is photoionizing the nebula and producing the Lyman continuum flux. The distances used for these calculations are listed in Table 1. Column 4 is the estimate mass of the spectral type of the ionizing star from Pecaut & Mamajek (2013).
@@ 4.2 Nature of the sources
By using previous studies on these sources, as well as results from our radio observations, we have assessed the nature of each individual source (see §3.1). From a sample point of view of the SOMA II sources, out of the 37 detections it is important to mention that three of them are well defined radio jets given their morphology and the estimated spectral index. For example, IRAS 16562-3959 and G339.88-1.26, are both very well previously studied ionized jets. Both have a central source that exhibits a jet-like morphology and a spectral index in the range for typical ionized jets associated with YSOs (Reynolds, 1986; Anglada et al., 1998; Tanaka et al., 2016), along with two additional sources classified as lobes that display non-thermal emission due to the synchrotron radiation. For G49.27-0.34 CM2A we report a spectral index in the typical range of ionized jets and two additional sources relatively close to the driving jet (CM2B at 0.7′′ and CM2C at 1.2′′), but both companions exhibit thermal emission.
We also propose a radio jet scenario for sources G35.58–0.03 and G305.20+0.21, based on the observed extended morphology, association with extended molecular outflows and evidence of knots from a non-thermal jet in the direction of the elongation. However, this classification remains less certain compared to other radio jets in the sample, and additional observations are needed to fully resolve and confirm the presence of the proposed non-thermal jets.
Additionally, we have three UC HII regions (G45.12+0.13 S14, G309.92+0.48 ATCA 1, and G35.58-0.03 31), following the suggestion of previous authors and the morphology of our detections and the reported spectral indices. The remaining sources are point sources being companion sources from the central detection or sources of an unknown nature, that could be of extragalactic nature or lower-mass YSOs. Most of them display a flat or negative spectral index suggesting non-thermal emission.
@@
In SOMA Radio I, we reported different evolutionary stages and nature of the sample, which included radio jets, HII regions, and variable stars. Four of the seven sources in that initial sample exhibited characteristics consistent with radio jets, such as elongation in the same direction as the associated molecular outflow with a central ionized jet and knots in aligned detections. In the sample presented in this paper, we observe a similar distribution: five of the nine regions show evidence of hosting a radio jet on the SOMA scale. Of these five, three (IRAS 16562-3959, G49.27-0.34, G339.88-1.26) display well-defined jets with knots, while the remaining two (G305B: G305.20+0.21 and G35.58-0.03) are proposed as jet candidates based on characteristics discussed in detail in Section 3, including spectral index, morphology, and multiplicity. Additionally, source G45.120.13, has been previously reported to be associated with an UCHII region which is consistent with our assessment from Figure 6 (see below). Furthermore, this source shows ionized emission at the lowest frequency (L-band) that resembles an hour-glass morphology similar to the very massive protostar G45.470.05 which is also associated with an UC HII region that is driving a powerful photoionized outflow and a jet candidate (Zhang et al., 2019a).
@@ 4.3 The Radio - Bolometric Luminosity Relation
Similar to Figure 3 in SOMA Radio I, in Figure 6 we present radio luminosity at 5 GHz from the Inner (left panel) and SOMA (right panel) scales versus bolometric luminosity. Note, when necessary, radio fluxes at 5 GHz have been estimated by scaling from adjacent frequencies using the derived spectral index. For the bolometric luminosity of SOMA sources (detections from SOMA I and II samples are squares and circles, respectively) we report the average of the “good” models from Table C1 of SOMA V and the error bar corresponds to the dispersion of these good models. We also show data for lower-mass YSOs associated with ionized jets from Anglada (1995) (small yellow dots). We scaled their fluxes from 3.6 cm, using a factor of 0.74, assuming that these sources have a spectral index 0.6, which is the expected value of ionized jets. A power-law fit to these data of is shown with a dashed line. UC/HC HII regions from Kurtz et al. (1994) are represented with a symbols. Note, the bolometric luminosities of the low-mass YSOs and the UC/HC HII regions are not measured in exactly the same way as the SED-fitting method of the SOMA sources, for which we are reporting intrinsic bolometric luminosities of the fitted SED models. A plot show the results based on isotropic bolometric luminosities is presented in Appendix A.
@@
We examine several theoretical models for the radio luminosity of massive protostars and show these in Figure 6. The black dotted line is the radio emission expected from an optically thin HII region, given the Lyman continuum luminosity of a single zero age main sequence (ZAMS) star at a given bolometric luminosity (Thompson, 1984). The light blue lines corresponds to the expected radio emission that arises from photoionization from a massive protostar forming via TCA, as predicted by the TTZ16 model, also for optically thin conditions at 5 GHz. These models have an initial core mass of and the three cases correspond to clump environment mass surface densities of and 3.16 g cm-2. The lower mass surface densities correspond to lower accretion rates, for which protostellar contraction towards the ZAMS occurs sooner, i.e., at lower masses and lower values of . Note, these models do not include any contribution from shock ionization.
@@
A model for estimating the radio emission from free-free emission from ionized gas produced by collisional ionization in shocks has been developed for a massive protostar forming from a core in a clump environment by Gardiner et al. (2024). This radio emission is also shown in Figure 6 with dark blue lines. We note that the contribution from shock ionization is quite spatially extended, with the emission from a 25,000 au radius aperture (more closely matching the SOMA scale) being about a factor of 10 higher than that from a 1,000 au aperture (more closely matching the inner scale).
@@
Comparing the distribution of SOMA Radio sources with the models, we note the following points. The SOMA Radio II sample, mostly being protostars of relatively high luminosity, help to better define the radio versus bolometric luminosity relation over the range from to , which is a regime where it is rising rapidly. We find that the radio luminosity measured on the SOMA scale rises steeply by about five dex as bolometric luminosity increases by 1 dex. Such a steep rise is consistent with basic expectations of TCA massive protostar models.
@@
Nevertheless, overall from the SOMA protostar sample we see that sources have relatively faint radio luminosities compared to predictions of protostellar models. This suggests that the rise in ionizing luminosity may occur only at relatively higher protostellar masses and luminosities. Such a relatively late rise in ionizing luminosity could be due to higher protostellar accretion rates or reflect systematic errors in the current protostellar models and their HII region radio emission (e.g., assumption all stellar EUV photons contribute to ionization / neglect of dust absorption). In addition, we caution that, at least on the SOMA scale, there is the possibility that we are missing radio flux due to interferometric filtering. A comparison with additional sets of theoretical models based on optically thin radio emission from the ionizing photon output of the TCA protostellar evolution models of Zhang & Tan (2018) is shown in Appendix B.
@@
We note that the HC/UC HII region sample of Kurtz et al. (1994) does include sources with higher radio luminosities, potentially indicating a systematic change arising from source evolution from the protostellar phase or else reflecting biases in sample selection from a single intrinsic protostellar population that has inherent large variation in radio luminosity.
@@
We see that the shock ionization model of Gardiner et al. (2024), which falls below the extrapolated lower-mass YSO relation, is consistent with the fainter end of the SOMA sample. Extending the comparison of the models with SOMA data on intermediate-mass protostars is a major goal of the SOMA Radio III paper (in prep.).

### v3

@@ 3.1.4 IRAS 16562-3959
Source 1 is the main detection in the region, and in our ATCA observations, this source is seen at all frequencies (5.5, 9, 17 and 22.8 GHz) and exhibits a morphology with a slight elongation in the NE-SW direction. SOMA II was able to resolve this source into two sources (G305B1 and G305B2) using high-spatial-resolution Gemini observations, finding morphologies that are also aligned in the NE-SW direction. They suggested that both sources are part of an outflow cavity. In our ATCA observations at the lowest frequencies (5.5 and 9 GHz), source 1 has two peaks, 1N and 1S (see inset in Figure 1), that appear to be associated with their MIR counterpart. Our global spectral index estimate of Source 1 is dominated by non-thermal emission ( -0.7), suggesting the existence of a high-speed jet in the region. Given the orientation of the sources, the driving source of this jet is likely source 1N, which is the brightest source, and we estimate a rough spectral index of .
@@
Source 4 is a faint compact eastern source only seen at the lowest frequencies (5.5 and 9 GHz), and we report a limit on its spectral index (see Table 7). On the other hand, source 3 is seen at all frequencies and presents an elongated morphology in the SE-NW direction that appears as a double peak at the longest contour levels shown in Figure 1. This source may be part of the HII region in A. J. Walsh et al. (1999), but in our observations we report it as an independent source. We report a negative spectral index ( -0.6) suggesting non-thermal emission.
Since Sources 2 and 4 are aligned in the same SW-NW direction, exhibit relatively compact morphologies and negative spectral indices, we conclude that the most likely nature of these sources is that they are non-thermal knots, centered in Source 1N. Following our convention for sources with jet-like morphology, we measured the flux densities at an intermediate scale that encompasses the central source and the jet knots (Sources 1, 2, and 4). At this scale, we estimate an upper limit on the spectral index of , which is indicative of a non-thermal jet in the region.
@@
| CM1 | 19:23:06.87 | +14:20:17.3 | 0.8(0.1) | IR | No |
| CM2A | 19:23:06.66 | +14:20:11.6 | 1.6(0.2) | IR | Yesa |
| CM2B | 19:23:06.62 | +14:20:12.1 | 0.7 | IR | Yesa |
| CM2C | 19:23:06.58 | +14:20:11.9 | 0.0(0.1) | IR | Yesa |
| CM3 | 19:23:06.71 | +14:19:59.0 | 0.5 | IR | Yesb |
@@
C. J. Cyganowski et al. (2011) suggests that CM1 is a single ionizing star of spectral type B0V based on optically thin free–free emission and the ionizing photon flux. Our 1.3 cm flux density values are one order of magnitude less than that reported by C. J. Cyganowski et al. (2011); A. P. M. Towner et al. (2017) with VLA observations and resolutions of ′′ (2.63 and 1.62 mJy). This difference is likely due to the different resolutions.
CM2: The MIR peak observed in SOMA II is associated with source CM2, previously reported as an unresolved, 4 detection at 0.027 mJy/beam rms at 3.6 cm (C. J. Cyganowski et al., 2011). Additionally, C. J. Cyganowski et al. (2011) estimated a spectral index limit of 0.2 at a resolution of ′′. Using our VLA data, we are able to resolve source CM2 into three components, which we have labeled as CM2A, CM2B and CM2C, with all sources detected at both 1.3 and 6 cm observations.
@@
CM3: From the contour maps in C. J. Cyganowski et al. (2011); A. P. M. Towner et al. (2017), source CM3 shows MIPS 24 m emission and SOMA II also detected this source in the SOFIA IR maps, but not in the NIR. This source has no previous radio detections at 1.3 and 3.6 cm. In our observations we are able to detect this source, but only at 6 cm (C-band).
CM3 exhibits a compact morphology and we report a relative flat spectral index with an upper limit of 0.5, indicative of thermal emission. To our knowledge, we are detecting this source for the first time in radio frequencies and its naming follows the convention of the main region, G49.27-0.34. This source may be a YSO at a more evolved stage than CM2, although additional observations are required for a definitive assessment.
@@
G339.88-1.26 (or IRAS 16484-4603) is a massive protostar located at 2.1 kpc (V. Krishnan et al., 2015), that has been previously observed at a variety of wavelengths. J. M. De Buizer et al. (2002) resolved the central MIR emission into three peaks (1A, 1B and 1C) and radio continuum observations have identified the central source is associated with a radio jet and the other two sources are associated with radio outflow lobes (S. P. Ellingsen et al., 1996; S. J. D. Purser et al., 2016), label as C, NE and SW. Using ALMA data, Y. Zhang et al. (2019a) reported two molecular outflows towards this region positioned at the source C. One of this outflows is located in the northeast-southwest orientation and exhibits the same angle as the ionized outflow reported by S. J. D. Purser et al. (2016).
@@
For our analysis, we used the ATCA observations at 5.5, 9.0, 17.0, and 22.8 GHz reported by S. J. D. Purser et al. (2016). We focus on the three sources within the SOMA scale i.e., NE, C, and SW, identified by S. J. D. Purser et al. (2016). Source C appears compact at higher frequencies but shows a slight extension at lower frequencies, particularly at 5.5 GHz. In contrast, Sources NE and SW exhibit relatively compact morphologies, with Source SW being especially compact.
As mentioned by S. J. D. Purser et al. (2016), source G339.88-1.26 shows characteristics consistent with an ionized radio jet driven by the central source C and two lobes of non-thermal emission. Given this jet-like morphology of the region we also measure the flux densities using an intermediate scale that encompass the three sources. We report a spectral index of 0.1 for the intermediate scale suggesting that the non-thermal emission from the lobes is more prevalent at this scale.
@@ 3.2 Radio SEDs
Building upon the results of SOMA II, in Figure 5 we present the E-SEDs, i.e., radio + IR SEDs, for our nine protostellar sources. The dashed lines correspond to the best fit to the radio continuum data using a power law of the form , where is the spectral index at the various scales: ”SOMA”; “Intermediate”; and “Inner”, as described in section 3. The spectral index was calculated using the flux density at the central frequencies from the images, so is calculated over a wide frequency range (20 GHz). The uncertainty in the spectral index was calculated with a Monte Carlo simulation that bootstrapped the flux density uncertainties. We estimated an upper limit in the spectral index for non-detections at higher frequencies using a value of of 3.
It is important to mention that at higher frequencies the free-free emission is likely contaminated by dust emission (see C. L. Brogan et al. (2016)), therefore adding some additional systematic uncertainty to the measurement of the free-free fluxes.
@@ 4.1 Multiplicity and new detections
To make sure we are not missing any UCHII regions, we estimate the physical properties from the 5 GHz sensitivity limit (rms) using equation 1 and 3 from S. Kurtz et al. (1994) and equation A.2.3 from N. Panagia & C. M. Walmsley (1978). These equations assume spherical symmetry and optically thin emission from a uniform-density plasma with . The results are listed in Table 9, where column 1 is the region name, column 2 is the logarithm of the Lyman continuum flux () required for ionization. With the and the results from Table II in N. Panagia (1973), we estimate the spectral type of the ionizing star (listed in column 3), further assuming that a single ZAMS star is photoionizing the nebula and producing the Lyman continuum flux. The distances used for these calculations are listed in Table 1. Column 4 is the estimate mass of the spectral type of the ionizing star from M. J. Pecaut & E. E. Mamajek (2013).
@@ 4.2 Nature of the sources
By using previous studies on these sources, as well as results from our radio observations, we have assessed the nature of each individual source (see §3.1). From a sample point of view of the SOMA II sources, out of the 37 detections it is important to mention that three of them are well defined radio jets given their morphology and the estimated spectral index. For example, IRAS 16562-3959 and G339.88-1.26, are both very well previously studied ionized jets. Both have a central source that exhibits a jet-like morphology and a spectral index in the range for typical ionized jets associated with YSOs (S. P. Reynolds, 1986; G. Anglada et al., 1998; K. E. I. Tanaka et al., 2016), along with two additional sources classified as lobes that display non-thermal emission due to the synchrotron radiation. For G49.27-0.34 CM2A we report a spectral index in the typical range of ionized jets and two additional sources relatively close to the driving jet (CM2B at 0.7′′ and CM2C at 1.2′′), but both companions exhibit thermal emission.
We also propose a radio jet scenario for sources G35.58–0.03, G305.21+0.21 (G305A) and G305.20+0.21 (G305B), based on the calculated spectral index, observed extended morphology, association with extended molecular outflows and in some cases evidence of knots from a non-thermal jet in the direction of the elongation. However, this classification remains less certain compared to other radio jets in the sample, and additional observations are needed to fully resolve and confirm the presence of the proposed non-thermal jets.
Additionally, we have three UC HII regions (G45.12+0.13 S14, G35.58-0.03 31 and IRAS 16562-3959 2), following the suggestion of previous authors and the morphology of our detections and the reported spectral indices. The remaining sources are point sources being companion sources from the central detection or sources of an unknown nature, that could be of extragalactic nature or lower-mass YSOs. Most of them display a flat or negative spectral index suggesting non-thermal emission. Table 10 summarizes all the sources reported in this paper along with their suggested nature. Column 1 lists the region, Column 2 the corresponding source name, and Column 3 the suggested nature of the source.
| Region | Source | Nature |
| G45.12+0.13 | S14 | UC HII Region |
| G45.12+0.13 | S20 | Extragalactic |
| G309.92+0.48 | ATCA1 | Optically Thick emitter |
| G309.92+0.48 | ATCA2 | Undetermined |
| G35.58-0.03 | 30A | Undetermined |
| G35.58-0.03 | 30B | Jet Knot (Candidate) |
| G35.58-0.03 | 30C | Jet Knot (Candidate) |
| G35.58-0.03 | 30D | Jet Knot (Candidate) |
| G35.58-0.03 | 30E | Jet Knot (Candidate) |
| G35.58-0.03 | 30F | Jet Knot (Candidate) |
| G35.58-0.03 | 30G | Jet Knot (Candidate) |
| G35.58-0.03 | 30H | Jet Knot (Candidate) |
| G35.58-0.03 | 30J | Jet Knot (Candidate) |
| G35.58-0.03 | 30K | Jet Knot (Candidate) |
| G35.58-0.03 | 30N | Unresolved Jet Candidate |
| G35.58-0.03 | 30S | Unresolved Jet Candidate |
| G35.58-0.03 | 31 | UC HII Region |
| IRAS 16562-3959 | C | Ionized Radio Jet |
| IRAS 16562-3959 | I-E | Non-thermal Lobe |
| IRAS 16562-3959 | I-W | Non-thermal Lobe |
| IRAS 16562-3959 | 1 | Extragalactic |
| IRAS 16562-3959 | 2 | HC HII Region |
| IRAS 16562-3959 | 3 | Low-mass YSO |
| IRAS 16562-3959 | 4 | Undetermined |
| G305.21+0.21 (G305A) | ATCA-1 | Jet Candidate |
| G305.20+0.21 (G305B) | ATCA-1 | Jet Candidate |
| G305.20+0.21 (G305B) | ATCA-2 | Jet Knot (Candidate) |
| G305.20+0.21 (G305B) | ATCA-3 | Part of the HII Region |
| G305.20+0.21 (G305B) | ATCA-4 | Jet Knot (Candidate) |
| G49.27-0.34 | CM1 | B0V Ionizing Star |
| G49.27-0.34 | CM2A | Jet Candidate |
| G49.27-0.34 | CM2B | Jet Knot (Candidate) |
| G49.27-0.34 | CM2C | Jet Knot (Candidate) |
| G49.27-0.34 | CM3 | Undetermined |
| G339.88-1.26 | C | Ionized Radio Jet |
| G339.88-1.26 | NE | Non-thermal Lobe |
| G339.88-1.26 | SW | Non-thermal Lobe |
@@
In SOMA Radio I, we reported different evolutionary stages and nature of the sample, which included radio jets, HII regions, and variable stars. Four of the seven sources in that initial sample exhibited characteristics consistent with radio jets, such as elongation in the same direction as the associated molecular outflow with a central ionized jet and knots in aligned detections. In the sample presented in this paper, we observe a similar distribution: six of the nine regions show evidence of hosting a radio jet on the SOMA scale. Of these six, three (IRAS 16562-3959, G49.27-0.34, G339.88-1.26) display well-defined jets with knots, while the remaining two (G305A: G305.21+0.21, G305B: G305.20+0.21 and G35.58-0.03) are proposed as jet candidates based on characteristics discussed in detail in Section 3, including spectral index, morphology, and multiplicity. Additionally, source G45.120.13, has been previously reported to be associated with an UCHII region which is consistent with our assessment from Figure 6 (see below). Furthermore, this source shows ionized emission at the lowest frequency (L-band) that resembles an hour-glass morphology similar to the very massive protostar G45.470.05 which is also associated with an UC HII region that is driving a powerful photoionized outflow and a jet candidate (Y. Zhang et al., 2019b).
@@ 4.3 The Radio - Bolometric Luminosity Relation
Similar to Figure 3 in SOMA Radio I, in Figure 6 we present radio luminosity at 5 GHz from the Inner (left panel) and SOMA (right panel) scales versus bolometric luminosity. Note, when necessary, radio fluxes at 5 GHz have been estimated by scaling from adjacent frequencies using the derived spectral index. For the bolometric luminosity of SOMA sources (detections from SOMA I and II samples are squares and circles, respectively) we report the average of the “good” models from Table C1 of SOMA V and the error bar corresponds to the dispersion of these good models. We also show data for lower-mass YSOs associated with ionized jets from G. Anglada (1995) (small yellow dots). We scaled their fluxes from 3.6 cm, using a factor of 0.74, assuming that these sources have a spectral index 0.6, which is the expected value of ionized jets. A power-law fit to these data of is shown with a dashed line. UC/HC HII regions from S. Kurtz et al. (1994) are represented with a symbols. Note, the bolometric luminosities of the low-mass YSOs and the UC/HC HII regions are not measured in exactly the same way as the SED-fitting method of the SOMA sources, for which we are reporting intrinsic bolometric luminosities of the fitted SED models. A plot show the results based on isotropic bolometric luminosities is presented in Appendix A.
@@
We examine several theoretical models for the radio luminosity of massive protostars and show these in Figure 6. The black dotted line is the radio emission expected from an optically thin HII region, given the Lyman continuum luminosity of a single zero age main sequence (ZAMS) star at a given bolometric luminosity (R. I. Thompson, 1984). The light blue lines corresponds to the expected radio emission that arises from photoionization from a massive protostar forming via TCA, as predicted by the TTZ16 model, also for optically thin conditions at 5 GHz. These models have an initial core mass of and the three cases correspond to clump environment mass surface densities of and 3.16 g cm-2. The lower mass surface densities correspond to lower accretion rates, for which protostellar contraction towards the ZAMS occurs sooner, i.e., at lower masses and lower values of . Note, these models do not include any contribution from shock ionization.
@@
A model for estimating the radio emission from free-free emission from ionized gas produced by collisional ionization in shocks has been developed for a massive protostar forming from a core in a clump environment by E. C. Gardiner et al. (2024). This radio emission is also shown in Figure 6 with dark blue lines. We note that the contribution from shock ionization is quite spatially extended, with the emission from a 25,000 au radius aperture (more closely matching the SOMA scale) being about a factor of 10 higher than that from a 1,000 au aperture (more closely matching the inner scale).
@@
Comparing the distribution of SOMA Radio sources with the models, we note the following points. The SOMA Radio II sample, mostly being protostars of relatively high luminosity, help to better define the radio versus bolometric luminosity relation over the range from to , which is a regime where it is rising rapidly. We find that the radio luminosity measured on the SOMA scale rises steeply by about four dex as bolometric luminosity increases by dex from to . Such a steep rise is consistent with basic expectations of TCA massive protostar models.
@@
Nevertheless, overall from the SOMA protostar sample we see that sources have relatively faint radio luminosities compared to predictions of protostellar models. This suggests that the rise in ionizing luminosity may occur only at relatively higher protostellar masses and luminosities. Such a relatively late rise in ionizing luminosity could be due to higher protostellar accretion rates or reflect systematic errors in the current protostellar models and their HII region radio emission (e.g., assumption all stellar EUV photons contribute to ionization / neglect of dust absorption). In addition, we caution that, at least on the SOMA scale, there is the possibility that we are missing radio flux due to interferometric filtering. A comparison with additional sets of theoretical models based on optically thin radio emission from the ionizing photon output of the TCA protostellar evolution models of Y. Zhang & J. C. Tan (2018) is shown in Appendix B.
The differences between the results shown in the two plots of Figure 6 (as well as Figures A1, B1, and B2), particularly for the sources we classified as having jet-like morphologies (G305.20+0.21, G305.21+0.21, IRAS 16562-3959, G339.881.26, and G49.270.34), arise from the different scales used to determine the radio luminosity (Ü. Kavak et al., 2021). As noted by W. O. Obonyo et al. (2024), when measuring well-resolved cores of radio jets, the sources tend to align with the fit of lower-mass YSOs associated with ionized jets (dashed line in Figure 6). For our inner scale (left plot), in most cases we were able to resolve the driving source of the radio jet and therefore all of our jet candidates fall in the low-mass fit within uncertainties, except for source G49.27-0.34.
The SOMA scale (right plot) takes into account all the emission in the region, adding contaminating emission from the jet knots and other sources within this region, further enhancing the radio luminosity. The artificial emission added by other sources and the different spatial scales, when compared to the inner scale, make the SOMA scale not as reliable to study the nature of the sources as the inner scale, but the SOMA scale still holds importance since it allows direct comparison with theoretical models with bigger scales that take into account jet knots and other multiple stellar sources in proto-clusters.
@@
We note that the HC/UC HII region sample of S. Kurtz et al. (1994) does include sources with higher radio luminosities, potentially indicating a systematic change arising from source evolution from the protostellar phase or else reflecting biases in sample selection from a single intrinsic protostellar population that has inherent large variation in radio luminosity.
@@
We see that the shock ionization model of E. C. Gardiner et al. (2024), which falls below the extrapolated lower-mass YSO relation, is consistent with the fainter end of the SOMA sample. Extending the comparison of the models with SOMA data on intermediate-mass protostars is a major goal of the SOMA Radio III paper (in prep.).